FIRST HIGH RESOLUTION ANALYSIS OF THE ν_{1}, ν_{2}, ν_{3}, ν_{4}, ν_{5}, AND ν_{6} BANDS OF COF^{35}Cl IN THE 600-2000 cm ^{−1}REGION

A. PERRIN, \underline{J.-M. FLAUD}, Laboratoire de Photophysique Moléculaire, CNRS, Université Paris Sud, Campus d’Orsay, Bat 350, 91405 Orsay Cedex, France; H. BÜRGER, G.PAWELKE, FB 9 - Anorganische Chemie Universität , D-42097 Wuppertal, Germany; S. SANDER, H.WILLNER, FB 6 - Anorganische Chemie Universität , D-47048 Duisburg, Germany.

Concern about the implication of chlorofluorocarbons (CFCs) on the amounts of ozone in the earth’s atmosphere has led to the measurement of the concentration of halogen containing compounds in the stratosphere. Of the various possible oxidation products of CFC’s in the stratosphere, carbonyl chlorofluoride COFCl may receive attention. COFCl is formed by the reaction of freon11 (CFCl_{3}) with O(^{1}D) oxygen atoms which are formed by photolysis of ozone ^{a}, ^{b}. High resolution atmospheric infrared spectra are potential means of monitoring COFCl. However, there is only little information on the spectroscopy of this molecule in the literature. We have recorded spectra of a ^{35}Cl enriched sample of COFCl at ca. 0.002 cm^{−1} resolution with the BRUKER Fourier transform spectrometer of the University of Wuppertal. We present here the first high resolution analyses of the six fundamentals ν_{1}, ν_{2}, ν_{3}, ν_{4}, ν_{5} and ν_{6} bands located at 1875.8, 1095.1, 764.4, 501.6, 408.8 and 666.6 cm^{−1} respectively for the ^{35}Cl isotopic species of COFCl. According to the results of the analyses, the ν_{3}, ν_{4}, ν_{5}, ν_{6} are unperturbed, while the analyses of the ν_{2} and ν_{1} bands was complicated by numerous resonances. Excellent results were obtained for the ν_{3}, ν_{4}, ν_{5}, ν_{6} bands and for the perturbed ν_{2} band. On the other hand, for the ν_{1} band for which the resonance scheme is more complex, the results of the analysis are still less satisfactory.

aJ.A.Kaye, A.R.Douglass, Ch.H.Jackman, R.S.Stolarski, R.Zander and G.Roland, J. Geophys. Res., \underline{D96}, 12865, (1991)

bF.Wu, and R.W.Carr, J. Phys. Chem. \underline{96}, 1743, (1992)